### v1

# Model-Based Reward Shaping for Adversarial Inverse Reinforcement Learning in Stochastic Environments

Abstract
In this paper, we aim to tackle the limitation of the Adversarial Inverse Reinforcement Learning (AIRL) method in stochastic environments where theoretical results cannot hold and performance is degraded. To address this issue, we propose a novel method which infuses the dynamics information into the reward shaping with the theoretical guarantee for the induced optimal policy in the stochastic environments. Incorporating our novel model-enhanced rewards, we present a novel Model-Enhanced AIRL framework, which integrates transition model estimation directly into reward shaping. Furthermore, we provide a comprehensive theoretical analysis of the reward error bound and performance difference bound for our method. The experimental results in MuJoCo benchmarks show that our method can achieve superior performance in stochastic environments and competitive performance in deterministic environments, with significant improvement in sample efficiency, compared to existing baselines.
1 Introduction
Reinforcement learning (RL) has achieved considerable success across various domains, including board game (Schrittwieser et al., 2020), MOBA game (Berner et al., 2019), time-delayed system (Wu et al., ; Wu et al., 2024), and cyber-physical systems (Wang et al., 2023a, b, c; Zhan et al., 2023). Despite these advances, RL highly depends on the quality of reward function design which demands expertise, intensive labour, and a great amount of time (Russell, 1998). To address this, imitation learning (IL) methods, such as Behavior Cloning (BC) (Torabi et al., 2018a) and Inverse Reinforcement Learning (IRL) (Arora & Doshi, 2021), leverage human or expert demonstrations to bypass the need for explicit reward functions. These methods aim to learn from the demonstrations to eventually match the distribution of expert behavior, and have shown great promise in applications like autonomous driving (Codevilla et al., 2018; Sun et al., 2018), legged locomotion (Peng et al., 2020; Ratliff et al., 2009), and planning tasks (Choudhury et al., 2018; Yin et al., 2022).
The notable approaches within IRL are Adversarial Imitation Learning (AIL) methods that build upon maximum entropy framework (Ziebart et al., 2008). These adversarial methods frame imitation learning as a maximum likelihood estimation problem on trajectory distributions, converting the distribution into a Boltzmann distribution parameterized by rewards under deterministic environment settings (Wu et al., 2024). This closely mirrors the distribution approximation found in generative models (Finn et al., 2016a; Swamy et al., 2021). Thus, model-free AIL approaches often follow generative model structures, such as GANs (Ho & Ermon, 2016; Fu et al., 2017) or diffusion models (Reuss et al., 2023), and require extensive sampling for distribution matching gradient updates (Orsini et al., 2021) conducting. Model-based IRL and AIL frameworks have also emerged, where model-based framework is designed to provide estimation for gradient and planning, leading to innovative combinations such as gradient-based IRL with model predictive control (MPC) (Das et al., 2021) and end-to-end differentiable IRL frameworks for complex robotics tasks (Baram et al., 2016, 2017; Sun et al., 2021; Rafailov et al., 2021). However, these approaches primarily address deterministic settings and struggle when applied to stochastic environments.
The only learning "deterministic" reward techniques among the existing AIL methods face significant performance degeneration in stochastic environments, leading to risk-seeking behavior and increased data requirements (Ziebart et al., 2010). For example, an agent trained under the deterministic Markov Decision Process (MDP) might aim to imitate expert behavior by seeking high rewards, yet fail to account for the low probability of some transitions in stochastic MDP settings. This happens because, in stochastic environments, the assumption that trajectory distributions are aligned with a Boltzmann distribution solely parameterized by deterministic rewards no longer holds. Instead, the dynamics information must also be incorporated into the formulation. There are two possible solutions. One is massive sampling to cover all possible outcomes, which is computationally expensive in large state action spaces (Devlin & Kudenko, 2011; Gupta et al., 2022). The other is estimating the dynamics information and integrating it into the reward design, making the reward "stochastic". Traditional reward design is usually based on state only $R(s_{t})$ (Torabi et al., 2018b), state-action pair $R(s_{t},a_{t})$ (Blondé & Kalousis, 2019), or transition tuple $R(s_{t},a_{t},s_{t+1})$ (Fu et al., 2017), where the information inputted can be thought as a deterministic sample piece under the stochastic setting. The challenge in stochastic environments calls for a different perspective of rewards – stochastic rewards absorbing the transition information.
Inspired by this idea, we propose a novel model-enhanced AIRL framework with a specifically tailored model-enhanced reward design approach to elevate performance in stochastic environments while remaining competitive in deterministic settings. In contrast to existing methods, our approach leverages the predictive power of the estimated transition model to shape rewards, represented as $\hat{R}(s_{t},a_{t},\hat{\mathcal{T}})$. This also enables us to generate synthetic trajectories to help guide policy optimization and reduce dependency on costly real-world interactions. As part of our analysis, we provide a theoretical guarantee on the optimal behavior for policies induced by our reward shaping and derive a bound on the performance gap with respect to the transition model errors. Empirically, we demonstrate that this integration significantly enhances sample complexity and policy performance in both settings, providing a comprehensive solution to the limitations of existing AIL methods in uncertain environments.
Contributions of this work include:
  • A novel reward shaping method with model estimation under the stochastic MDP setting, which provides the optimal policy invariance guarantee.
  • A novel model-based adversarial IRL framework that seamlessly incorporates transition model training, adversarial reward learning with model estimation and forward model-based RL process, enhancing performance in stochastic environments, and sample efficiency.
  • Theoretical analysis on reward learning with model estimation under the adversarial framework and performance difference under transition model learning errors.
  • Empirical validation that demonstrates our approach’s performance improvements in stochastic environments and comparable performance in deterministic environments as well as significant improvement in sample efficiency.
In Sec. 2, we introduce related works in AIL and reward shaping. In Sec. 3, we provide the necessary preliminaries for MDP and IRL. In Sec. 4, we present our model-enhanced reward shaping method and corresponding theoretical guarantee. In Sec. 5, we present our Model-Enhanced AIRL framework design together with theoretical analysis on reward error bound and performance difference bound. In Sec. 6, we show the experimental results in Mujoco for various benchmarks. Sec. 7 concludes the paper.
2 Related Works
Adversarial Imitation Learning.
Margin optimization based IRL methods (Ng et al., 2000; Abbeel & Ng, 2004; Ratliff et al., 2006) aim to learn reward functions that explain expert behavior better than other policies by a margin. Bayesian approaches were introduced with different prior assumptions on reward distributions, such as Boltzmann distributions (Ramachandran & Amir, 2007; Choi & Kim, 2011; Chan & van der Schaar, 2021) or Gaussian Processes (Levine et al., 2011). Other statistical learning methods include multi-class classification (Klein et al., 2012; Brown et al., 2019) and regression trees (Levine et al., 2010). The entropy optimization approach has seen significant development. To avoid biases from maximum margin methods, the maximum entropy principle (Shore & Johnson, 1980) is used to infer distributions over trajectories parameterized by reward weights. Ziebart et al. (2008, 2010) proposed a Lagrangian dual framework to cast the reward learning into a maximum likelihood problem with linear-weighted feature-based reward representation. Wulfmeier et al. (2015) extended the framework to nonlinear reward representations, and Finn et al. (2016b) combines importance sampling techniques to enable model-free estimation. Inspired by GANs, adversarial methods were introduced for policy and reward learning in IRL (Ho & Ermon, 2016; Fu et al., 2017; Torabi et al., 2018b). However, these methods typically work with Maximum Entropy (ME) formulation yet suffer from sample inefficiency. Although there have been efforts to combine adversarial methods with off-policy RL agents to improve sample efficiency (Kostrikov et al., 2018; Blondé & Kalousis, 2019; Blondé et al., 2022), few extend it to the model-based setting which might further the improvement, and none of these approaches addresses the rewards learning in stochastic MDP settings.
Rewards Shaping.
Reward shaping (Dorigo & Colombetti, 1994; Randløv & Alstrøm, 1998) is a technique that enhances the original reward signal by adding additional domain information, making it easier for the agent to learn optimal behavior. This can be defines as $\hat{R}=R+F$, where $F$ is the shaping function and $\hat{R}$ is the shaped reward function. Potential-based reward shaping (PBRS) (Ng et al., 2000) builds the potential function on states, $F(s,a,s^{\prime})=\phi(s^{\prime})-\phi(s)$, while ensuring the policy invariance property, which refers to inducing the same optimal behavior under different rewards $R$ and $\hat{R}$. Nonetheless, there also exist other variants on the inputs of the potential functions such as state-action (Wiewiora et al., 2003), state-time (Devlin & Kudenko, 2012), and value function (Harutyunyan et al., 2015) as potential function input. There are also some latest attempts of reward shaping without utilization of domain knowledge potential function to solve exploration under sparse rewards (Hu et al., 2020; Devidze et al., 2022; Gupta et al., 2022; Skalse et al., 2023).
MBIRL.
Model-Based RL (MBRL) has emerged as a promising direction for improving sample efficiency and generalization (Janner et al., 2019; Yu et al., 2020). MBRL combines various learned dynamics neural network structures with planning (Hansen et al., 2022; Sikchi et al., 2022).This framework has been successfully extended to vision-based control tasks (Hafner et al., 2019; Zhan et al., 2023). Integrating IRL with MBRL has also shown success. For example, Das et al. (2021) and Herman et al. (2016) presented a gradient-based IRL approach using different policy optimization methods with dynamic models for linear-weighted features reward learning. In Das et al. (2021), the dynamic model is used to pass forward/backward the gradient in order to update the IRL and policy optimization modules. Similarly, end-to-end differentiable adversarial IRL frameworks to various state spaces have also been explored (Baram et al., 2016, 2017; Sun et al., 2021; Rafailov et al., 2021), where dynamic model serves a similar role. Despite these advancements, existing methods rarely address the specific challenges posed by stochastic environments, which limit reward learning performance. To our knowledge, this is the first study that provide a theoretical analysis on the performance difference with learned dynamic model for the adversarial IRL problem under stochastic MDP.
3 Preliminaries
MDP.
RL is usually formulated as a Markov Decision Process (MDP) $\mathcal{M}$ (Puterman, 2014) denoted as a tuple $\langle\mathcal{S},\mathcal{A},\mathcal{T},\gamma,R,\rho_{0}\rangle$. $\rho_{0}$ is the initial distribution of the state. $s\in\mathcal{S},a\in\mathcal{A}$ stands for the state and action space respectively. $\mathcal{T}$ stands for the transition dynamic such that $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$. $\gamma\in(0,1)$ is the discounted factor, $R$ stands for reward function such that $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ and $\|R\|_{\infty}\leq R_{\max}$. The discounted visitation distribution of trajectory $\tau$ with policy $\pi$ is given by:
$p(\tau)=\rho_{0}\prod_{t=0}^{T-1}\gamma^{t}\mathcal{T}(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t}).$ (1)
The objective function of RL is $\max\mathbb{E}_{\tau\sim p(\tau)}\left[\sum_{t=0}^{T}\gamma^{t}R(\tau)-H(\pi)\right]$, where $H$ is the log likelihood of the policy. We introduce Soft Value Iteration for bellmen update (Haarnoja et al., 2018), where $Q^{soft}$ and $V^{soft}$ denotes the soft Q function and Value function respectively:
$V^{soft}(s_{t})$ $=log\sum_{a_{t}\in\mathcal{A}}\exp{Q^{soft}(s_{t},a_{t})}da_{t},$ (2)
$Q^{soft}(s_{t},a_{t})$ $=R(s_{t},a_{t})+\gamma\mathbb{E}_{\mathcal{T}}\left[V^{soft}(s_{t+1})|s_{t},a_{t}\right],$ (3)
$\pi(a_{t}|s_{t})$ $=\exp{(Q^{soft}(s_{t},a_{t})-V^{soft}(s_{t}))},$ (4)
where the soft Advantage function is defined as $A^{soft}(s_{t},a_{t})=Q^{soft}(s_{t},a_{t})-V^{soft}(s_{t})$.
Inverse RL.
In IRL setting, we usually consider the MDP without reward as $\mathcal{M}^{\prime}$ where $R$ is also unknown. We denote the data buffer $\mathcal{D}_{exp}$ which collects trajectories from an expert policy $\pi^{E}$. We consider a reward function $R_{\theta}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, where $\theta$ is the reward parameter. An IRL problem can be defined as a pair $\mathcal{B}=(\mathcal{M}^{\prime},\pi^{E})$. A reward function $R_{\theta}$ is feasible for $\mathcal{B}$ if $\pi^{E}$ is an optimal policy for the MDP $\mathcal{M}^{\prime}\cup R_{\theta}$, and we denote the set of feasible rewards as $\mathcal{R}_{\mathcal{B}}$. Using maximize likelihood estimation framework, we can formulate the IRL as the following maximum causal entropy problem:
$\arg\max_{\theta}\mathbb{E}_{\tau\sim\mathcal{D}_{exp}}\log p_{\theta}(\tau),$ (5)
where $Q_{R_{\theta}}^{soft}$ and $V^{soft}_{R_{\theta}}$ are based on $R_{\theta}$ and $p_{\theta}(\tau)\propto\rho_{0}\prod_{t=0}^{T-1}\mathcal{T}(s_{t+1}|s_{t},a_{t})\exp(Q^{soft}_{\hat{R}_{\theta}}(s_{t},a_{t})-V^{soft}_{R_{\theta}}(s_{t}))$ (Ziebart et al., 2010). Under deterministic MDP, the above problem can be simplified as maximum entropy, where $p_{\theta}(\tau)\propto\frac{1}{Z_{\theta}}\exp{\sum_{t=0}^{T-1}R_{\theta}(s_{t},a_{t})}$ and $Z_{\theta}$ is the temperature factor of the Boltzmann Distribution (Ziebart et al., 2008).
4 Model Estimation in reward shaping
Table 1: We summarize the different reward formulations and their dynamic properties in this table. Components refer to the input pair that the reward functions take. Reward Shaping indicates whether there is the extra physical potential information involved where X means no reward shaping used. Dynamics information shows whether transitions are involved in the reward function.
Methods Components Reward Shaping Dynamics Information
AIRL (Fu et al., 2017) $s_{t},a_{t},s_{t+1}$ $R(s_{t},a_{t})+\gamma\phi(s_{t+1})-\phi(s_{t})$ single sample
AIRL(State Only) $s_{t}$ $R(s_{t})+\textrm{constant}$ X
DAC (Kostrikov et al., 2018) $s_{t},a_{t}$ X X
SAM (Blondé & Kalousis, 2019) $s_{t},a_{t}$ X X
SQIL (Reddy et al., 2019) $s_{t},a_{t}$ binary X
GAIfO (Torabi et al., 2018b) $s_{t},s_{t+1}$ X single sample
Ours $s_{t},a_{t},\mathcal{T}$ $R(s_{t},a_{t})+\gamma\mathbb{E}_{\mathcal{T}}[\phi(s_{t+1})|s_{t},a_{t}]-\phi(s_{t})$ transition model
In this section, we illustrate the advantages of involving transition dynamics into the reward shaping, especially in stochastic MDP settings. Most of literature work has various formulations and definitions (Table 1), but few considers transition dynamic information in the reward shaping. Defining rewards solely based on states, $R^{s}(s_{t})$, offers limited utility in environments where actions are critical. Even though the state-action pair-based rewards $R^{sa}(s_{t},a_{t})$ can capture the missing information of the taken action, it fails to consider any future information, the successive state $s_{t+1}$. Transition tuple-based rewards $R^{tuple}(s_{t},a_{t},s_{t+1})$ incorporate the dynamics information in a sampling-based way, which requires abundant data to learn the underlying relationship of two consecutive states, potentially raising the sample efficiency issue in the stochastic environment with the huge state space. To address this issue, we propose dynamics-based rewards shaping $\hat{R}(s_{t},a_{t},\mathcal{T})$, which explicitly infuse the dynamics information $\mathcal{T}$ on the potential function, thus significantly improving the sample-efficiency. Specifically, our rewards shaping is defined as
$\hat{R}(s_{t},a_{t},\mathcal{T})=R(s_{t},a_{t})+\gamma\mathbb{E}_{\mathcal{T}}\left[\phi(s_{t+1})|s_{t},a_{t}\right]-\phi(s_{t}),$ (6)
where $\phi$ is a state-only potential function, $\mathcal{T}$ is the dynamics. Another insight of the above reward shaping is to resemble the advantage function with the soft value function as the potential function, which we will elaborate in Sec. 5.1. With the given reward shaping $\hat{R}$, it is crucial to show that it induces the same optimal behaviour as the ground-true reward $R$. We formally define this policy invariance property as follows.
Definition 4.1.
(Memarian et al., 2021) Let $R$ and $\hat{R}$ be two reward functions. We say they induce the same soft optimal policy under transition dynamics $\mathcal{T}$ if, for all states $s\in\mathcal{S}$ and actions $a\in\mathcal{A}$:
$A_{R}^{soft}(s_{t},a_{t})=A_{\hat{R}}^{soft}(s_{t},a_{t}).$ (7)
With the above definition, we can transfer the proof of policy invariant property of our designed reward shaping (Eq. 6) to showing the equivalence of soft advantage functions, which is proved in the following theorem. The detailed proof can be found in A.1.
Theorem 4.2 (Reward Invariance).
Let $R$ and $\hat{R}$ be two reward functions. $R$ and $\hat{R}$ induce the same soft optimal policy under all transition dynamics $\mathcal{T}$ if $\hat{R}(s_{t},a_{t},\mathcal{T})=R(s_{t},a_{t})+\gamma\mathbb{E}_{\mathcal{T}}[\phi(s_{t+1})|s_{t},a_{t}]-\phi(s_{t})$ for some potential-shaping function $\phi:\mathcal{S}\rightarrow\mathbb{R}$.
Thm. 4.2 implies that the optimal policy induced from our model-enhanced rewards shaping $\hat{R}$ (Eq. 6) is equivalent to the optimal policy trained by the ground-truth reward function $R$ under the soft Value Iteration fashion.
5 Model Enhanced Adversarial IRL
In this section, we first elaborate on the adversarial formulation of our reward shaping (Eq. 8) and present the theoretical insight (Proposition 5.1) of the equivalence between cross-entropy training loss of adversarial reward shaping formulation and maximum log-likelihood loss of original maximum causal entropy IRL problem. Then, in the Sec. 5.2, we showcase our practical algorithm framework with trajectory generation and transition model learning in the loop, as shown in  Fig. 1. Furthermore, we theoretically investigate the reward function bound (Thm. 5.3) and performance difference bound (Thm. 5.4) under the transition model learning error.
Figure 1: Framework overview of Model-Enhanced Adversarial IRL. Different color arrows stand for different sample flows. Purple stands for real environmental interaction samples, pink stands for synthetic samples generated from learned transition model, and blue stands for mixed of both.
5.1 Adversarial Formulation of Reward Shaping
In order to learn a robust reward function, we apply the AIRL framework (Fu et al., 2017). Inspired by GANs (Goodfellow et al., 2014), the idea behind AIRL is to train a binary discriminator $D(s_{t},a_{t},s_{t+1})$ to distinguish state-action-transition samples from an expert and those generated by imitator policy. However, as mentioned above, we only take in state-action pair and transition function to define our reward function which also extends to our discriminator as follows:
$D_{\theta}(s_{t},a_{t},\mathcal{T})=\frac{\exp\{f_{\theta}(s_{t},a_{t},\mathcal{T})\}}{\exp\{f_{\theta}(s_{t},a_{t},\mathcal{T})\}+\pi(a_{t}|s_{t})},$ (8)
where $f_{\theta}(s_{t},a_{t},\mathcal{T})=R_{\theta}(s_{t},a_{t})+\gamma\mathbb{E}_{\mathcal{T}}\left[\phi_{\theta}(s_{t+1})|s_{t},a_{t}\right]-\phi_{\theta}(s_{t})$ resembles the reward shaping defined above. The loss function for the training discriminator is defined below.
$\mathcal{L}_{disc}=-\mathbb{E}_{\mathcal{D}_{exp}}\left[\log D_{\theta}(s,a,\mathcal{T})\right]-\mathbb{E}_{\pi}\left[\log(1-D_{\theta}(s,a,\mathcal{T}))\right].$ (9)
With the model-enhanced stochastic reward shaping, we can bridge this adversarial formulation with the original maximum causal entropy IRL problem. In the following proposition, we give a sketch of proof to show the link between the loss of discriminator and loss of MCE IRL, which sheds the light on how this adversarial formulation can recover the rewards of the task. Proof details can be found in Appendix B.1.
Proposition 5.1.
Consider an undiscounted MDP. Suppose $f_{\theta}$ and $\pi$ at the current iteration are the soft-optimal advantage function and policy for reward function $R_{\theta}$. Minimising the cross-entropy loss of the discriminator under generator $\pi$ is equivalent to maximising the log-likelihood under MCE IRL.
With above given proposition, we can construct a direct intuition that $f^{*}_{\theta}$ should be equal to $\hat{R}_{\theta}$ the reward shaping we introduced early and resemble the soft advantage function. To extract rewards to represent reward used for policy optimization, we use $\log(D_{\theta}(s,a,\mathcal{T}))-\log(1-D_{\theta}(s,a,\mathcal{T}))$, which resembles the entropy-regularized reward shaping $f_{\theta}(s,a,\mathcal{T})-\log\pi(a|s)$. Given this entropy-regularized reward, it is straightforward to see why the optimal policy can satisfy the RL objectives, $\max\mathbb{E}_{\tau\sim p(\tau)}\left[\sum_{t=0}^{T-1}\gamma^{t}\hat{R}_{\theta}(\tau)-H(\pi)\right]$.
5.2 Algorithm Framework
Algorithm 1 Model Enhanced Adversarial IRL
1:  Obtain expert buffer $\mathcal{D}_{exp}$.
2:  Initialize policy $\pi$, discriminator $D_{\theta}$, buffers $\mathcal{D}_{env},\mathcal{D}_{gen}$, and transition model $\hat{\mathcal{T}}$.
3:  for step $t$ in {1, …, N} do
4:     Interact with real environments and add state-action pair to $\mathcal{D}_{env}$.
[…]
6:        Pretrain transition model $\hat{\mathcal{T}}$.
7:     else
8:        Sample state-action batch from $\mathcal{D}_{exp},\mathcal{D}_{env}$ respectively.
9:        Train $D_{\theta}$ via cross entropy loss  Eq. 9 to classify expert data from samples.
10:        Update dynamic model $\hat{\mathcal{T}}$ with MLE loss and generate $H$-steps trajectories to $\mathcal{D}_{gen}$.
11:        Sample state-action batches from $\mathcal{D}_{env},\textrm{and }\mathcal{D}_{gen}$ with varying ratio.
12:        Update $\pi$ with respect to $\hat{R}_{\theta}$ using Soft Actor Critic policy optimization.
13:     end if
14:  end for
In this section, we present the overall framework of Model-Enhanced Adversarial IRL and illustrate how transition model training is incorporated into the learning loop. We assume the estimated transition distribution $\hat{\mathcal{T}}(\cdot|s,a)$ follows a Gaussian distribution with mean and standard deviation parameterized by the MLP, and the model is updated with standard maximum likelihood loss. At each iteration, the updated transition model is utilized for reward calculation and synthetic data generation, which is stored in the synthetic trajectory replay buffer. Unlike AIRL and GAIL, our framework operates in an off-policy manner, where samples used for both discriminator and policy update are drawn from a combination of the environmental replay buffer and the synthetic replay buffer. An overview of our framework is shown in  Fig. 1, and detailed algorithmic steps are provided in  Alg. 1 and Appendix D.
Sample Efficiency: To improve sample efficiency, we leverage the estimated transition model to generate $H$-steps synthetic trajectories data alongside real interaction data, facilitating policy optimization. Given that the estimated transition model is inaccurate at the beginning, we employ a dynamic ratio between real and synthetic data to prevent the model from being misled by unlikely synthetic transitions (Janner et al., 2019; Zhan et al., 2023). Specifically, early-stage generated trajectories are not stored persistently, unlike real interactions which are fully stored in the off-policy environmental replay buffer. To maintain training stability, we use a synthetic replay buffer with a size that gradually increases as training progresses, ensuring a balanced inclusion of synthetic data over time. The growth rates of the data ratio and buffer size are adjusted based on the complexity of the transition model learning process and can be fine-tuned via hyperparameters.
Distribution Shift: To mitigate distribution shift (Lee et al., 2020; Lin et al., 2020) during training, we employ a strategy involving the learned transition model. Typically, during interaction, the real state $s_{t}$ is used as input to the actor, and the resulting action $a_{t}$ is applied in the environment. To incorporate the transition model, we predict a synthetic state $\hat{s}_{t}$ from previous $s_{t-1}$ and $a_{t-1}$. This generated $\hat{s}_{t}$ is then fed into the actor to produce action $\hat{a}_{t}$. The actions $a_{t}$ and $\hat{a}_{t}$ are mixed and applied to the environment with a certain ratio, and the resulting pairs $(s_{t},a_{t})$ or $(s_{t},\hat{a}_{t})$ are stored in the environmental replay buffer. This approach helps balance the exploration of real and model-predicted dynamics, reducing the impact of distributional discrepancies.
5.3 Performance Analysis
[…]
Assumption 5.2 (Transition Model Error).
Since transition model is trained through a supervised fashion, we can use a PAC generalization bound (Shalev-Shwartz & Ben-David, 2014) for sample error. Therefore, we assume that the total variation distance between $\mathcal{T}$ and $\hat{\mathcal{T}}$ is bounded by $\epsilon_{\mathcal{T}}$ through $[0,T]$:
$\max_{t}\mathbb{E}_{s\sim\pi_{D,t}}\left[D_{TV}(\mathcal{T}(s^{\prime}|s,a)|\hat{\mathcal{T}}(s^{\prime}|s,a))\right]\leq\epsilon_{\mathcal{T}},$ (10)
which is a common assumption that adopted in literature (Janner et al., 2019; Sikchi et al., 2022).
Next, with the assumed total visitation bound on transition models ( 5.2), we can bound the error on rewards learning through model-enhanced reward shaping.
Theorem 5.3 (Reward Function Error Bound).
Let $\mathfrak{B}=(\mathcal{M}^{\prime},\pi^{*})$ and $\hat{\mathfrak{B}}=(\hat{\mathcal{M}^{\prime}},\pi^{*})$ be two IRL problems, then for any $R^{E}\in\mathcal{R}_{\mathfrak{B}}$ there is a corresponding $\hat{R}^{E}\in\mathcal{R}_{\hat{\mathfrak{B}}}$ such that
$\|R^{E}-\hat{R}^{E}\|\leq\frac{\gamma}{1-\gamma}|\mathcal{S}|\epsilon_{\mathcal{T}}R_{\max}.$ (11)
Proof of  Thm. 5.3 can be found in Appendix C.3. With rewards bound above, we can extend the bound to the value function, which represents the performance difference brought up by estimated transition model error under RL setting.
Theorem 5.4 (Performance Difference Bound).
The performance difference between the optimal policies ($\pi^{*}$ and $\hat{\pi}^{*}$) in corresponding MDPs ($\mathcal{M}^{\prime}\cup R$ and ^$\hat{\mathcal{M}}^{\prime}\cup\hat{R}$) can be bounded as follows:
$\|V_{\mathcal{M}^{\prime}\cup R}^{\pi^{*}}-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\hat{\pi}^{*}}\|\leq\epsilon_{\mathcal{T}}\left[\frac{\gamma}{(1-\gamma)^{2}}R_{\max}+\frac{1+\gamma}{(1-\gamma)^{2}}R_{\max}|\mathcal{S}|\right].$ (12)
The detailed proof of  Thm. 5.4 is presented in Appendix C.6. The above theorem highlights the relationship between the performance gap and the transition model error, also implying that a perfectly-learned transition model ($\epsilon_{\mathcal{T}}\rightarrow 0$) could make the performance difference negligible.
6 Experiments
In this section, we evaluate the performance and sample efficiency of our Model-Enhanced Adversarial IRL framework. We aim to demonstrate the superiority of our method in stochastic environments, achieving better performance and sample efficiency compared to existing approaches. Additionally, in deterministic settings, our method not only maintains competitive performance, but also achieves better sample efficiency with baselines. All experiments are conducted on the MuJoCo benchmarks (Todorov et al., 2012). To simulate stochastic dynamics in MuJoCo, we introduce the agent-unknown Gaussian noise with a mean of 0 and a standard deviation of 0.5 to the environmental interaction steps. All the expert data are generated by an expert trained with standard SAC under deterministic or stochastic MuJoCo environments. Our experiments are designed to highlight the key advantages of our framework:
  • Superiority in Stochastic Environments: In stochastic settings, our method significantly outperforms other approaches, consistently surpassing expert-level performance more rapidly. This enhanced ability to learn under uncertainty is attributed to our framework’s effectiveness in leveraging model-based synthetic data.
  • Performance in Deterministic Environments: We demonstrate that our method is either better or competitive with existing AIL methods’ performances in deterministic settings.
  • Sample Efficiency For both deterministic and stochastic settings, our method can reach expert performance with fewer training steps than all the other baselines.
We primarily compare our approach with other Adversarial Imitation Learning (AIL) methods, including the on-policy algorithms GAIL (Ho & Ermon, 2016) and AIRL (Fu et al., 2017), and the off-policy method Discriminator Actor-Critic (DAC) (Kostrikov et al., 2018). For policy optimization, we use Proximal Policy Optimization (PPO) (Schulman et al., 2017) for both GAIL and AIRL, and Soft Actor-Critic (SAC) (Haarnoja et al., 2018) for both DAC and our proposed method. All implementations of PPO and SAC are referenced from the Clean RL library (Huang et al., 2022). Each algorithm is trained with 100k environmental steps and evaluated each 1k steps across 5 different seeds for tasks including InvertedPendulum-v4 and InvertedDoublePendulum-v4. For Hopper-v3, AIRL and GAIL are trained with 10M steps and evaluated each 100k steps across 5 different seeds, but DAC and our algorithm are trained with 1M environmental steps and evaluated each 10k steps across 5 different seeds. We conduct the experiments under 3 different random seeds to seek more general trend and conclude systematic results. All the experiments are run on the Desktop equipped with RTX 4090 and Core-i9 13900K. The learning curves of all methods are provided in Appendix E.
Performance in Stochastic MuJoCo.
In  Table 2, we present the performance under the stochastic MuJoCo environments. Our method can achieve the best performance compared to all of the baselines in stochastic environments. Specifically, for InvertedDoublePendulum-v4, introducing stochasticity into the dynamics significantly degrades the performance. Our method still can still achieve performance better than all baselines. Furthermore, in the stochastic settings, the performance of DAC is decreased significantly, as DAC’s ineffective reward formulation on state-action pairs under uncertain conditions, also resulting a trend of learning instability (Appendix E). For Hopper-v4, our method remarkably surpasses the all baselines with significantly large margins. These results indicate that our approach can effectively learn the better reward function from demonstrations in stochastic environments, resulting in significant performance improvement.
Table 2: Performance of expert and all algorithms in stochastic MuJoCo Environments.
Environment Expert GAIL AIRL DAC Ours
InvertedPendulum-v4(Stoch) $1000.0_{\pm 0.0}$ $1000.0_{\pm 80.8}$ $1000.0_{\pm 25.4}$ $1000.0_{\pm 69.2}$ $1000.0_{\pm 31.2}$
InvertedDoublePendulum-v4(Stoch) $210.1_{\pm 98.6}$ $276.9_{\pm 25.1}$ $234.9_{\pm 38.1}$ $142.3_{\pm 31.2}$ $286.7_{\pm 39.2}$
Hopper-v3(Stoch) $2257.8_{\pm 317.4}$ $2832.0_{\pm 105.5}$ $2834.0_{\pm 333.2}$ $2408.0_{\pm 78.6}$ $3139.0_{\pm 109.9}$
Performance in Deterministic MuJoCo.
The performance of deterministic MuJoCo environments can be found in  Table 3. For the tasks with deterministic dynamics, our method can achieve the performance aligning with all of baselines and the expert in InvertedDoublePendulum-v4, InvertedDoublePendulum-v4 and Hopper-v3. Additionally, our method shows superiority in learning stability (Appendix E) and sample efficiency (Table 4). Therefore, our method shows competitive performance with the baselines in the deterministic environments with better sample efficiency and learning stability.
Table 3: Performance of expert and all algorithms in deterministic MuJoCo Environments.
Environment Expert GAIL AIRL DAC Ours
InvertedPendulum-v4 $1000.0_{\pm 0.0}$ $1000.0_{\pm 0.0}$ $1000.0_{\pm 0.0}$ $1000.0_{\pm 0.0}$ $1000.0_{\pm 0.0}$
InvertedDoublePendulum-v4 $9320.0_{\pm 578.8}$ $9322.0_{\pm 39.6}$ $9359.0_{\pm 88.3}$ $9360.0_{\pm 12.3}$ $9361.0_{\pm 0.2}$
Hopper-v3 $3329.3_{\pm 306.6}$ $3485.0_{\pm 5.6}$ $3472.0_{\pm 3.6}$ $3495.0_{\pm 4.5}$ $3463.0_{\pm 2.7}$
Sample Efficiency.
In  Table 4, we report the number of training steps required to achieve expert-level performance. Our method shows significant superiority in sample efficiency across all of the benchmarks under both deterministic and stochastic settings. Specifically, for InvertedPendulum-v4, all four methods are able to reach expert-level performance. However, our method reaches expert performance in the fewest training steps, $9.7k$ and $11.7k$ for deterministic and stochastic settings respectively. In the case of InvertedDoublePendulum-v4, introducing stochasticity into the dynamics makes it challenging for all algorithms to reach expert-level performance from noisy demonstrations. Especially, the strong baseline, DAC, failed to reach the expert’s performance, while our method can arrive at the expert-level performance within $27.3k$ steps. For Hopper-v3, all four methods achieve expert-level performance, but our method reaches it with the fewest samples whether in deterministic ($195.3k$) or stochastic ($170.0k$) settings.
Table 4: Averaged steps used for all algorithms to reach experts’ performances in deterministic and stochastic Environments. $X$ stands for that the algorithm does not reach the expert performance within given training steps.
Environment GAIL AIRL DAC Ours
InvertedPendulum-v4 $37.7k$ $38.0k$ $15.3k$ $9.7k$
InvertedPendulum-v4(Stoch) $54.0k$ $58.0k$ $11.7k$ $11.7k$
InvertedDoublePendulum-v4 $87.0k$ $91.5k$ $28.3k$ $11.3k$
InvertedDoublePendulum-v4(Stoch) $72.3k$ $75.3k$ X $27.3k$
Hopper-v3 $5400.0k$ $6400.0k$ $243.5k$ $195.3k$
Hopper-v3(Stoch) $4170.0k$ $3200.0k$ $786.7k$ $170.0k$
7 Conclusion
In this paper, we presented a novel model-enhanced adversarial inverse reinforcement learning framework by incorporating model-based techniques with reward shaping, specifically designed to enhance performance in stochastic environments and maintain competitive performance in deterministic setting while significantly improve performance under the traditional AIL approaches. The theoretical analysis provides guarantees on the optimal policy invariance under the transition model involved reward shaping and highlight the relationship between performance gap and transition model error, showing that the gaps becomes negligible with a well-learned model. Empirical evaluations on Mujoco benchmark environments validate the effectiveness of our method, showcasing its superior performance and sample efficiency across different tasks. Future works will focus on further refining the model estimation process to handle more complex and dynamic environments and exploring extensions of the framework to multi-agent and hierarchical reinforcement learning scenarios. Additionally, it would be valuable to investigate the generalization ability of our framework in a transfer learning tasks. Overall, our approach offers a promising direction for advancing model-based adversarial IRL, with the potential to scale to a broader range of real-world applications.
8 Reproducible Statement
This work uses the open-source MuJoCo (Todorov et al., 2012) as the benchmark. The practical implementation of our method is built on the CleanRL repository (Huang et al., 2022). All hyperparameters to reproduce our experimental results, including learning rates and transition model settings, are explicitly listed in Appendix D. For every reported result, we averaged the performance over three random seeds, and the seed initialization is included for exact reproducibility.
Appendix A Reward Shaping Soft Optimal Policy
Theorem A.1.
Let $R$ and $\hat{R}$ be two reward functions. $R$ and $\hat{R}$ induce the same soft optimal policy under all transition dynamics $\mathcal{T}$ if $\hat{R}(s_{t},a_{t})=R(s_{t},a_{t})+\gamma\mathbb{E}_{\mathcal{T}}[\phi(s_{t+1})|s_{t},a_{t}]-\phi(s_{t})$ for some potential-shaping function $\phi:\mathcal{S}\rightarrow\mathbb{R}$.
Proof.
According to Soft VI ( Eq. 2), we can expand the representation of $Q^{soft}_{\hat{R}}(s_{t},a_{t})$ as follows.
$Q^{soft}_{\hat{R}}(s_{t},a_{t})$ $=R(s_{t},a_{t})+\gamma\mathbb{E}_{\mathcal{T}}\left[\phi(s_{t+1})|s_{t},a_{t}\right]-\phi(s_{t})+\gamma\mathbb{E}_{\mathcal{T}}\left[V_{\hat{R}}^{soft}(s_{t+1})|s_{t},a_{t}\right],$
$Q^{soft}_{\hat{R}}(s_{t},a_{t})+\phi(s_{t})$ $=R(s_{t},a_{t})+\gamma\mathbb{E}_{\mathcal{T}}\left[V_{\hat{R}}^{soft}(s_{t+1})+\phi(s_{t+1})|s_{t},a_{t}\right],$
$Q^{soft}_{\hat{R}}(s_{t},a_{t})+\phi(s_{t})$ $=R(s_{t},a_{t})+\gamma\mathbb{E}_{\mathcal{T}}\left[\log\sum_{a\in\mathcal{A}}\exp\left(Q_{\hat{R}}^{soft}(s_{t+1},a)+\phi(s_{t+1})\right)|s_{t},a_{t}\right].$
From above induction, we can tell that $Q^{soft}_{\hat{R}}(s_{t},a_{t})+\phi(s_{t})$ satisfy the soft bellmen update with original $R$. Thus, with simple induction, we can arrive that $Q_{R}^{soft}(s_{t},a_{t})=Q_{\hat{R}}^{soft}(s_{t},a_{t})+\phi(s_{t})$. Then, we can derive the advantage function
$A_{\hat{R}}^{soft}(s_{t},a_{t})$ $=Q_{\hat{R}}^{soft}(s,a)-V_{\hat{R}}^{soft}(s_{t})$
$=Q_{\hat{R}}^{soft}(s_{t},a_{t})-\log\sum_{a\in\mathcal{A}}\exp\left(Q_{\hat{R}}^{soft}(s_{t},a_{t})\right)$
$=Q_{\hat{R}}^{soft}(s_{t},a_{t})+\phi(s_{t})-\log\sum_{a\in\mathcal{A}}\exp\left(Q_{\hat{R}}^{soft}(s_{t},a_{t})+\phi(s_{t})\right)$
$=Q_{R}^{soft}(s_{t},a_{t})-\log\sum_{a\in\mathcal{A}}\exp\left(Q_{R}^{soft}(s_{t},a_{t})\right)$
$=A_{R}^{soft}(s_{t},a_{t}).$
Appendix B Adversarial Reward Learning
Proposition B.1.
Consider an undiscounted MDP. Suppose $f_{\theta}$ and $\pi$ at current iteration are the soft-optimal advantage function and policy for reward function $R_{\theta}$. Minimising the cross-entropy loss of the discriminator under generator $\pi$ is equivalent to maximising the log-likelihood under MCE IRL.
Proof.
$\mathcal{L}_{IRL}(\mathcal{D}_{exp},\theta)$ $=\mathbb{E}_{\mathcal{D}_{exp}}[\log p_{\theta}(\tau)]$
$=\mathbb{E}_{\mathcal{D}_{exp}}\left[\sum_{t=0}^{T-1}\log\pi(a_{t}|s_{t})+\log\rho_{0}+\sum_{t=1}^{T}\mathcal{T}(s_{t+1}|s_{t},a_{t})\right]$
$=\mathbb{E}_{\mathcal{D}_{exp}}\left[\sum_{t=0}^{T-1}\left(Q^{soft}_{\theta}(s_{t},a_{t})-V^{soft}_{\theta}(s_{t})\right)\right]+\text{constant}.$
Breaking down above equations with soft VI (Eq. 2), we can arrive the following.
$\mathbb{E}_{\mathcal{D}_{exp}}\left[\sum_{t=0}^{T-1}R_{\theta}(s_{t},a_{t})\right]+\mathbb{E}_{\mathcal{D}_{exp}}\left[\sum_{t=0}^{T-2}\mathbb{E}_{\mathcal{T}}\left[V_{\theta}^{soft}(s_{t+1})|s_{t},a_{t}\right]\right]-\mathbb{E}_{\mathcal{D}_{exp}}\left[\sum_{t=0}^{T-1}V_{\theta}^{soft}(s_{t})\right].$ (13)
Next we will derive the gradient of the loss.
$\nabla_{\theta}\mathcal{L}(\mathcal{D}_{exp},\theta)=\underbrace{\nabla_{\theta}\mathbb{E}_{\mathcal{D}_{exp}}\left[\sum_{t=0}^{T-1}R_{\theta}(s_{t},a_{t})\right]}_{A}+\\ \underbrace{\nabla_{\theta}\mathbb{E}_{\mathcal{D}_{exp}}\left[\sum_{t=0}^{T-2}\left(\mathbb{E}_{\mathcal{T}}\left[V_{\theta}^{soft}(s_{t+1})|s_{t},a_{t}\right]\right)-V_{\theta}^{soft}(s_{t+1})\right]}_{B}-\underbrace{\nabla_{\theta}V_{\theta}^{soft}(s_{0})}_{C}.$ (14)
Let’s get explicit expression of each part.
$A$ $=\mathbb{E}_{\mathcal{D}_{exp}}\left[\sum_{t=0}^{T-1}\nabla_{\theta}R_{\theta}(s_{t},a_{t})\right]$
$C$ $=\nabla_{\theta}\log\sum_{a_{t}\in\mathcal{A}}\exp{Q_{\theta}^{soft}(s_{t},a_{t})}$
$=\sum_{a_{t}\in\mathcal{A}}\pi(a_{t}|s_{t})\nabla_{\theta}Q_{\theta}^{soft}(s_{t},a_{t})$
$=\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}\nabla_{\theta}R_{\theta}(s_{t},a_{t})\right].$
The transitions in the demonstration $\mathcal{D}_{exp}$ will follow $\mathcal{T}$. Thus we can drop the $\mathbb{E}_{\mathcal{T}}$, and $B$ term will cancel out, ending up to 00. To summarize, the gradient of log MLE loss of MCE IRL is the following.
$\nabla_{\theta}\mathcal{L}_{IRL}(\mathcal{D}_{exp},\theta)=\mathbb{E}_{\mathcal{D}_{exp}}\left[\sum_{t=0}^{T-1}\nabla_{\theta}R_{\theta}(s_{t},a_{t})\right]-\mathbb{E}_{\pi}\left[\sum_{t=0}^{T-1}\nabla_{\theta}R_{\theta}(s_{t},a_{t})\right].$ (15)
Next, we will start to derive the gradient of cross-entropy discriminator training loss. Remember the discriminator loss is defined in Eq 9.
$\log D_{\theta}(s,a)$ $=f_{\theta}(s,a)-\log(\exp\{f_{\theta}(s,a)\}+\pi(a|s)),$
$\log(1-D_{\theta}(s,a))$ $=\log\pi(a|s)-\log(\exp\{f_{\theta}(s,a)\}+\pi(a|s)).$
Then, the gradient of each term is as follow:
$\nabla_{\theta}\log D_{\theta}(s,a)$ $=\nabla_{\theta}f_{\theta}(s,a)-\frac{\exp\{f_{\theta}(s,a)\}\nabla_{\theta}f_{\theta}(s,a)}{\exp\{f_{\theta}(s,a)\}+\pi(a|s)},$
$\nabla_{\theta}\log(1-D_{\theta}(s,a))$ $=-\frac{\exp\{f_{\theta}(s,a)\}\nabla_{\theta}f_{\theta}(s,a)}{\exp\{f_{\theta}(s,a)\}+\pi(a|s)}.$
Since $\pi$ is trained by using $f_{\theta}$ as shaped reward, from soft VI we can derive that $\pi^{*}_{f_{\theta}}(a|s)=\exp A^{soft}_{f_{\theta}}(s,a)$. By assumption, we assume that $f_{\theta}$ is the advantage function of $R_{\theta}$, $f_{\theta}(s,a)=A^{soft}_{R_{\theta}}(s,a)$. From Thm 4.2, we know that $A^{soft}_{R_{\theta}}(s,a)=A^{soft}_{f_{\theta}}(s,a)$, which also implies that $\pi^{*}_{f_{\theta}}=\pi^{*}_{R_{\theta}}$. Then, we can deduce the gradient of the loss of discriminator.
$-\nabla_{\theta}\mathcal{L}_{disc}$ $=\mathbb{E}_{\mathcal{D}_{exp}}\left[\nabla_{\theta}\log D_{\theta}(s,a)\right]+\mathbb{E}_{\pi}\left[\nabla_{\theta}\log(1-D_{\theta}(s,a))\right]$
$=\mathbb{E}_{\mathcal{D}_{exp}}\left[\frac{1}{2}\nabla_{\theta}f_{\theta}(s,a)\right]-\mathbb{E}_{\pi}\left[\frac{1}{2}\nabla_{\theta}f_{\theta}(s,a)\right],$
$-2\nabla_{\theta}\mathcal{L}_{disc}$ $=\mathbb{E}_{\mathcal{D}_{exp}}\left[\nabla_{\theta}f_{\theta}(s,a)\right]-\mathbb{E}_{\pi}\left[\nabla_{\theta}f_{\theta}(s,a)\right].$
Appendix C Performance Gap Analysis
Lemma C.1 (Implicit Feasible Reward Set (Ng et al., 2000)).
Let $\mathfrak{B}=(\mathcal{M}^{\prime},\pi^{*})$ be an IRL problem. Then $R\in\mathcal{R}_{\mathfrak{B}}$ if and only if for all $(s,a)\in\mathcal{S}\times\mathcal{A}$ the following holds:
$Q_{\mathcal{M}^{\prime}\cup R}^{\pi^{*}}(s,a)-V_{\mathcal{M}^{\prime}\cup R}^{\pi^{*}}(s)$ $=0\quad\textrm{if}\;\pi^{*}(a|s)>0,$
$Q_{\mathcal{M}^{\prime}\cup R}^{\pi^{*}}(s,a)-V_{\mathcal{M}^{\prime}\cup R}^{\pi^{*}}(s)$ $\leq 0\quad\textrm{if}\;\pi^{*}(a|s)=0.$
Combined with the traditional Value Iteration of RL, we can write out the explicit form of the reward function $R$.
Lemma C.2 (Explicit Feasible Reward Function (Metelli et al., 2021)).
With the above lemma conditions, $R\in\mathcal{R}_{\mathfrak{B}}$ if and only if there exist $\xi\in\mathbb{R}_{\geq 0}^{\mathcal{S}\times\mathcal{A}}$ and value function $V\in\mathbb{R}^{\mathcal{S}}$ such that:
$R(s,a)=V(s)-\gamma\sum_{s^{\prime}\in\mathcal{S}}\mathcal{T}(s^{\prime}|s,a)V(s^{\prime})-\xi(s,a)\mathbb{I}\{\pi^{*}(a|s)=0\}.$ (16)
With  Eq. 16, we can derive the following error bound between $R\in\mathcal{R}^{E}_{\mathfrak{B}}$ and $\hat{R}^{E}\in\mathcal{R}_{\hat{\mathfrak{B}}}$.
Theorem C.3 (Reward Function Error Bound).
Let $\mathfrak{B}=(\mathcal{M}^{\prime},\pi^{*})$ and $\hat{\mathfrak{B}}=(\hat{\mathcal{M}^{\prime}},\pi^{*})$ be two IRL problems, then for any $R^{E}\in\mathcal{R}_{\mathfrak{B}}$ there is a corresponding $\hat{R}^{E}\in\mathcal{R}_{\hat{\mathfrak{B}}}$ such that
$\|R^{E}-\hat{R}^{E}\|\leq\frac{\gamma}{1-\gamma}|\mathcal{S}|\epsilon_{\mathcal{T}}R_{\max}.$ (17)
Proof.
From  Lem. C.2, we can derive the following representations of $R$ and $\hat{R}$ with the same set of $V$ and $\xi$:
$R^{E}(s,a)$ $=V(s)-\gamma\sum_{s^{\prime}\in\mathcal{S}}\mathcal{T}(s^{\prime}|s,a)V(s^{\prime})-\xi(s,a)\mathbb{I}\{\pi^{*}(a|s)=0\},$
$\hat{R}^{E}(s,a)$ $=V(s)-\gamma\sum_{s^{\prime}\in\mathcal{S}}\hat{\mathcal{T}}(s^{\prime}|s,a)V(s^{\prime})-\xi(s,a)\mathbb{I}\{\pi^{*}(a|s)=0\}.$
The difference between $R^{E}$ and $\hat{R}^{E}$ can be bounded as follows:
$\|R^{E}-\hat{R}^{E}\|$ $\leq\gamma\sum_{s^{\prime}\in\mathcal{S}}D_{TV}\left(\mathcal{T}(s^{\prime}|s,a)|\hat{\mathcal{T}}(s^{\prime}|s,a)\right)\cdot\|V(s^{\prime})\|.$
Given that the total variation distance between the two dynamics is bounded by $\epsilon_{\mathcal{T}}$, and the reward function is bounded by $R_{\max}$, together with the definition of the value function, we have $\|V\|_{\infty}\leq\frac{R_{\max}}{1-\gamma}$. Substituting these bounds, we derive the following inequality:
$\|R^{E}-\hat{R}^{E}\|\leq\frac{\gamma}{1-\gamma}|\mathcal{S}|\epsilon_{\mathcal{T}}R_{\max}.$
Next, we will propagate this bound to the value functions of optimal policy regarding different reward functions $R$ and $\hat{R}$. From the traditional Value iteration, we can write out the value function.
$V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s)=\sum_{a\in\mathcal{A}}\pi(a|s)\sum_{s^{\prime}\in\mathcal{S}}\mathcal{T}(s^{\prime}|s,a)\left[R(s,a)+\gamma V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s^{\prime})\right].$ (18)
Lemma C.4 (Value Function Error under same policy and different rewards and MDP).
$||V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s)-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi}(s)||$: the performance difference of the same policy in different MDPs.
$||V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s)-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi}(s)||\leq\epsilon_{\mathcal{T}}\frac{1+\gamma}{(1-\gamma)^{2}}R_{\max}|\mathcal{S}|.$ (19)
Proof.
$||V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s)-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi}(s)||$ (20)
$\leq\sum_{a\in\mathcal{A}}\pi(a|s)\sum_{s^{\prime}\in\mathcal{S}}||\mathcal{T}(s^{\prime}|s,a)\left[R(s,a)+\gamma V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s^{\prime})\right]-\hat{\mathcal{T}}(s^{\prime}|s,a)\left[R(s,a)+\gamma V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s^{\prime})\right]$
$+\hat{\mathcal{T}}(s^{\prime}|s,a)\left[R(s,a)+\gamma V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s^{\prime})\right]-\hat{\mathcal{T}}(s^{\prime}|s,a)\left[\hat{R}(s,a)+\gamma V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi}(s^{\prime})\right]||$
$\leq\sum_{a\in\mathcal{A}}\pi(a|s)\sum_{s^{\prime}\in\mathcal{S}}(\epsilon_{\mathcal{T}}\frac{R_{\max}}{1-\gamma}+\hat{\mathcal{T}}(s^{\prime}|s,a)\epsilon_{\mathcal{T}}(\frac{\gamma R_{\max}}{1-\gamma}|\mathcal{S}|+\gamma||V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s^{\prime})-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi}(s^{\prime})||))$
$=\sum_{a\in\mathcal{A}}\pi(a|s)(\epsilon_{\mathcal{T}}\frac{R_{\max}}{1-\gamma}|\mathcal{S}|+\epsilon_{\mathcal{T}}\frac{\gamma R_{\max}}{1-\gamma}|\mathcal{S}|+\gamma||V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s^{\prime})-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi}(s^{\prime})||)$
$\leq\epsilon_{\mathcal{T}}\frac{1+\gamma}{1-\gamma}R_{\max}|\mathcal{S}|+\gamma||V_{\mathcal{M}^{\prime}\cup R}^{\pi}(s^{\prime})-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi}(s^{\prime})||$
$\leq\epsilon_{\mathcal{T}}\frac{1+\gamma}{(1-\gamma)^{2}}R_{\max}|\mathcal{S}|.$
Lemma C.5.
Let $\|V_{\hat{\mathcal{M}^{\prime}}\cup\hat{R}}^{\pi_{1}}(s)-V_{\hat{\mathcal{M}^{\prime}}\cup\hat{R}}^{\pi_{2}}(s)\|$ denote the performance difference between different policies $\pi_{1}$ and $\pi_{2}$ in the same learned MDP (Viano et al., 2021; Zhang et al., 2020). The following inequality holds:
$\|V_{\hat{\mathcal{M}^{\prime}}\cup\hat{R}}^{\pi_{1}}(s)-V_{\hat{\mathcal{M}^{\prime}}\cup\hat{R}}^{\pi_{2}}(s)\|\leq\frac{\gamma}{(1-\gamma)^{2}}\epsilon_{\mathcal{T}}R_{\max}.$
Theorem C.6 (Performance Difference Bound).
The performance difference between the optimal policies ($\pi^{*}$ and $\hat{\pi}^{*}$) in corresponding MDPs ($\mathcal{M}^{\prime}\cup R$ and ^$\hat{\mathcal{M}}^{\prime}\cup\hat{R}$) can be bounded as follows:
$\|V_{\mathcal{M}^{\prime}\cup R}^{\pi^{*}}-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\hat{\pi}^{*}}\|\leq\epsilon_{\mathcal{T}}\left[\frac{\gamma}{(1-\gamma)^{2}}R_{\max}+\frac{1+\gamma}{(1-\gamma)^{2}}R_{\max}|\mathcal{S}|\right].$ (21)
Proof.
$||V_{\mathcal{M}^{\prime}\cup R}^{\pi^{*}}(s)-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\hat{\pi}^{*}}(s)||$
$\leq||V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\pi^{*}}(s)-V_{\hat{\mathcal{M}^{\prime}}\cup\hat{R}}^{\hat{\pi}^{*}}(s)||+||V_{\mathcal{M}^{\prime}\cup R}^{\hat{\pi}^{*}}(s)-V_{\hat{\mathcal{M}}^{\prime}\cup\hat{R}}^{\hat{\pi}^{*}}(s)||$
$=\epsilon_{\mathcal{T}}\frac{\gamma}{(1-\gamma)^{2}}R_{\max}+\epsilon_{\mathcal{T}}\frac{1+\gamma}{(1-\gamma)^{2}}R_{\max}|\mathcal{S}|$
$=\epsilon_{\mathcal{T}}\left[\frac{\gamma}{(1-\gamma)^{2}}R_{\max}+\frac{1+\gamma}{(1-\gamma)^{2}}R_{\max}|\mathcal{S}|\right].$
Appendix D Implementation Details
We used two identical 2-layer MLP with 100 hidden units and ReLU activiations for the reward term $r$ and the shaping term $\phi$. The neural nets are trained using Adam optimizer (Kingma, 2014) with learning rate being $0.0003$. For the learned dynamic model, we also used a 2-layer MLP with ReLU activations and 128 hidden units. Gradient clipping is added on the model network with the value of 10. The model network is trained by the Adam optimizer as well with learning rate $0.0003$. For SAC and PPO, we follow the implementation from CleanRL repository (Huang et al., 2022).
Appendix E Graph Results
Below is the testing return diagrams from both deterministic and stochastic Mujoco Environments
Figure 2: Training return diagram of benchmarks and our algorithm in the InvertedPendulum-v4. We only show the training return until reaching the experts performance of all the other benchmarks, but the thorough training return of our algorithm.
Figure 3: Training return diagram of benchmarks and our algorithm in the InvertedDoublePendulum-v4.
Figure 4: Training return diagram of benchmarks and our algorithm in the Hopper-v3.